OpenStudy (anonymous):

How did you get -6? x^2+4x-12=0

OpenStudy (phi):

do you know how to factor this ?

OpenStudy (phi):

if you need more info on factoring, this might help (when you have time) https://www.khanacademy.org/math/algebra/multiplying-factoring-expression/factoring-quadratic-expressions/v/factoring-quadratic-expressions

OpenStudy (phi):

but to answer your question x^2+4x-12 can be factored into (x+6)(x-2) if you multiply that out you will get x^2 +6x -2x -12 which simplifies to x^2 +4x -12 (i.e. what we started with) \[ x^2+4x-12=0 \\ \text{ becomes } \\(x+6)(x-2) = 0 \] when you multiply two "things" and get zero, that must mean one or the other (or both) is zero so either x+6 =0 or x-2 = 0 the first equation x+6=0 gives us x+6 -6 = 0 -6 or x = -6 the second equation gives us x=2 so there are two different x values that make your equation true.

OpenStudy (anonymous):

yes i have for i have. x^2+4x+4=12 +4 then i ended up with (x+2)^2= 16 which gave me x +/- 4 x+2=4 = 2 x-2= -4 = -2 but i type the equation into my calculator (y=), i'm supposed to end up with x=2 and x= -6

OpenStudy (phi):

(x+2)^2= 16 take the square root of both sides \[ x+2 = \pm 4 \] we should write that as two separate equations: x+2 = +4 and x+2 = -4 (notice this is different from what you wrote: x-2 = -4. )

OpenStudy (anonymous):

thank you so much. (:

OpenStudy (phi):

yw